I. Solve polynomial equation using formula method



3x² +9x + 7=0




Ii. Solve polynomial equation using factorization method



X³+2x² ‐ 9x -18=0





1
Expert's answer
2021-12-06T16:02:25-0500

i.


"3x\u00b2 +9x + 7=0"

"D=(9)^2-4(3)(7)=-3"

"x=\dfrac{-9\pm\sqrt{-3}}{2(3)}=-\dfrac{3}{2}\pm\dfrac{\sqrt{3}}{6}i"

"\{-\dfrac{3}{2}-\dfrac{\sqrt{3}}{6}i,-\dfrac{3}{2}+\dfrac{\sqrt{3}}{6}i\}"



ii.

"x^3+2x^2 -9x-18=0"

"(x^3+2x^2)-(9x+18)=0"

"x^2(x+2)-9(x+2)=0"

"(x+2)(x^2-9)=0"

"(x+2)(x+3)(x-3)=0"

"x_1=-2, x_2=-3, x_3=3"

"\{-3, -2, 3\}"